### v1

# A general comparison principle for the pluripotential complex Monge-Ampère flow

Abstract.
We prove a comparison principle for the pluripotential complex Monge-Ampère flows for the right-hand side of the form $dt\wedge d\mu$ where $d\mu$ is dominated by a Monge-Ampère measure of a bounded plurisubharmonic function. As a consequence, we obtain the uniqueness of the weak solution to the pluripotential Cauchy-Dirichlet problem. We also study the long-term behavior of the solution under some assumption.
1. Introduction
In this paper, we aim at extending parabolic pluripotential theory introduced by Guedj, Lu and Zeriahi in [GLZ21-2]. Pluripotential theory for the complex Monge-Ampère equation has been well developed, starting from the works by Bedford and Taylor [BT76, BT82]. Guedj, Lu and Zeriahi [GLZ20, GLZ21-2] successfully developed the parabolic pluripotential theory and studied the weak solutions for the pluripotential complex Monge-Ampère flows. They proved the existence and uniqueness of the solution to the Cauchy-Dirichlet problem when the measure on the right-hand side is given a positive $L^{p}$ density function for some $p>1$. We considerably extend their results by studying the Cauchy-Dirichlet problem for general measures.
We recall the setup of [GLZ21-2], which will be used throughout this paper. Let $\Omega\Subset\mathbb{C}^{n}$ be a bounded strictly pseudoconvex domain. From now on, we fix $0<T<+\infty$ and denote $\Omega_{T}=(0,T)\times\Omega$. We say that a function $u:\Omega_{T}\rightarrow[-\infty,+\infty)$ is locally uniformly Lipschitz in $(0,T)$ if for any compact $J\Subset(0,T)$, there exists $\kappa_{J}>0$ satisfying
$u(t,z)\leq u(s,z)+\kappa_{J}\lvert t-s\rvert\text{ for all $t,s\in J$ and for all $z\in\Omega$.}$
Next, we say that a function $u:\Omega_{T}\rightarrow\mathbb{R}$ is locally uniformly semi-concave in $(0,T)$ if for any compact $J\Subset(0,T)$, there exists $C_{J}>0$ such that
$t\mapsto u(t,z)-C_{J}t^{2}\text{ is concave in $J$ for all $z\in\Omega$.}$
The family of parabolic potentials $\mathcal{P}(\Omega_{T})$ is the set of functions $u$ such that
  • for any $t\in(0,T)$, $u(t,\cdot)\in PSH(\Omega)$,
  • $u$ is locally uniformly Lipschitz in $(0,T)$.
The Cauchy-Dirichlet boundary data $h$ is a function defined on $\partial_{0}\Omega_{T}:=([0,T)\times\partial\Omega)\cup(\{0\}\times\Omega)$ satisfying
  • the restriction of $h$ on $0,T)\times\partial\Omega$ is continuous,
  • the family $\{h(\cdot,z)~:~z\in\partial\Omega\}$ is locally uniformly Lipschitz in $(0,T)$,
  • for any $\zeta\in\partial\Omega$,
    $h_{0}:=h(0,\cdot)\in PSH(\Omega)\cap L^{\infty}(\Omega)\text{ and }\lim_{\Omega\ni z\rightarrow\zeta}h(0,z)=h(0,\zeta).$
Let $F(t,z,r):[0,T)\times\Omega\times\mathbb{R}\rightarrow\mathbb{R}$ be a continuous function which is increasing in $r$, bounded in $0,T)\times\Omega\times J$ for each $J\Subset\mathbb{R}$ and $(t,r)\mapsto F(t,\cdot,r)$ is locally uniformly Lipschitz and semi-convex. Let $\mu$ be a non-negative Borel measure on $\Omega$ such that $PSH(\Omega)\subset L^{1}_{loc}(\Omega,\mu)$.
We consider the following complex Monge-Ampère flows
(1.1) $dt\wedge(dd^{c}u)^{n}=e^{\partial_{t}u+F(t,z,u)}dt\wedge d\mu\text{ in }\Omega_{T},$
where $u\in\mathcal{P}(\Omega_{T})\cap L^{\infty}(\Omega_{T})$ is an unknown function. Here, $dt\wedge(dd^{c}u)^{n}$ is understood in the pluripotential sense as defined in [GLZ21-2, Definition 2.2]. The Cauchy-Dirichlet problem for the (1.1) with the Cauchy-Dirichlet boundary data $h$ is finding $u\in\mathcal{P}(\Omega_{T})\cap L^{\infty}(\Omega_{T})$ such that $u$ solves (1.1) and
(1.2) $\begin{cases}&\lim_{(t,z)\rightarrow(\tau,\zeta)}u(t,z)=h(\tau,\zeta)\text{ for all }(\tau,\zeta)\in[0,T)\times\partial\Omega,\\ &\lim_{t\rightarrow 0+}u(t,\cdot)=h_{0}\text{ in }L^{1}(\Omega,d\mu).\end{cases}$
Let $dV$ be a Lebesgue measure on $\mathbb{C}^{n}$. Guedj, Lu and Zeriahi [GLZ21-2] proved the existence and uniqueness of the solution to the Cauchy-Dirichlet problem with a measure $\mu$ and a boundary data $h$ satisfying
$d\mu=gdV\text{ for some }g\in L^{p}(\Omega),~p>1\text{ with }g>0\text{ a.e.},$
and there exists $C>0$ such that
(1.3) $t\lvert\partial_{t}h(t,z)\rvert\leq C\text{ for all }(t,z)\in(0,T)\times\partial\Omega,$
and
[…]
We focus on their uniqueness result. The authors first proved the following comparison principle and then used this to prove the uniqueness of the solution which is locally uniformly semi-concave in $(0,T)$.
Theorem 1.1.
[GLZ21-2, Theorem 6.6] Let $d\mu=gdV$ for some $g\in L^{p}(\Omega)$ with $p>1$ and $g>0$ a.e. with respect to the Lebesgue measure on $\Omega$. Let $u,v\in\mathcal{P}(\Omega_{T})\cap L^{\infty}(\Omega_{T})$ with the Cauchy-Dirichlet boundary data $h_{1}$, $h_{2}$ as in (1.2). Assume that
  1. (a)
    $v$ is locally uniformly semi-concave in $t\in(0,T)$,
  2. (b)
    $dt\wedge(dd^{c}u)^{n}\geq e^{\partial_{t}u+F(t,z,u)}dt\wedge d\mu$ in $\Omega_{T}$,
  3. (c)
    $dt\wedge(dd^{c}v)^{n}\leq e^{\partial_{t}v+F(t,z,v)}dt\wedge d\mu$ in $\Omega_{T}$,
  4. (d)
    $h_{1}$ satisfies (1.3) and (1.4).
If $h_{1}\leq h_{2}$, then $u\leq v$.
Our main result generalizes this comparison principle for a measure $\mu$ such that there exists $\varphi\in PSH(\Omega)\cap L^{\infty}(\Omega)$ satisfying
(1.5) $\begin{cases}&(dd^{c}\varphi)^{n}\geq d\mu\text{ in }\Omega,\\ &\lim_{z\rightarrow\partial\Omega}\varphi(z)=0.\end{cases}$
Thanks to the Kołodziej’s subsolution theorem, it is equivalent to the existence of $\psi\in PSH(\Omega)\cap L^{\infty}(\Omega)$ satisfying
(1.6) $\begin{cases}&(dd^{c}\psi)^{n}=d\mu\text{ in }\Omega,\\ &\lim_{z\rightarrow\partial\Omega}\psi(z)=0.\end{cases}$
From now, we fix such $\mu$, $\varphi$ and $\psi$ throughout this paper.
Theorem 1.2.
Let $u,v\in\mathcal{P}(\Omega_{T})\cap L^{\infty}(\Omega_{T})$ with the Cauchy-Dirichlet boundary data $h_{1}$, $h_{2}$ as in (1.2). Assume that
  1. (a)
    $u$ and $v$ are locally uniformly semi-concave in $(0,T)$,
  2. (b)
    $dt\wedge(dd^{c}u)^{n}\geq e^{\partial_{t}u+F(t,z,u)}dt\wedge d\mu$ in $\Omega_{T}$,
  3. (c)
    $dt\wedge(dd^{c}v)^{n}\leq e^{\partial_{t}v+F(t,z,v)}dt\wedge d\mu$ in $\Omega_{T}$,
  4. (d)
    $h_{1}$ satisfies (1.3).
If $h_{1}\leq h_{2}$, then $u\leq v$.
Remark 1.3.
Note that the conditions (a) and (d) in Theorem 1.1 are different from the ones of Theorem 1.2. In [GLZ21-2], instead of comparing a subsolution and a supersolution, the authors compared an upper envelope of a subsolution and a supersolution. They used the assumption (1.4) in [GLZ21-2, Theorem 4.8] to show that the upper envelope is locally uniformly semi-concave, which was used to prove the (joint) continuity of the envelope [GLZ21-2, Theorem 6.4, Theorem 6.5]. Also, due to their assumption on the measure, they could obtain the continuity of a supersolution. In this paper, we do not have the (joint) continuity of the upper envelope and a supersolution in general, so we need a different approach. Instead, we directly compare two locally uniformly semi-concave subsolution and supersolution without the assumption (1.4).
The existence of a solution to the Cauchy-Dirichlet problem under assumptions (1.3), (1.4), and (1.5) is proved in our previous work [Ka25]. By using Theorem 1.2, we now obtain the uniqueness of the solution to the Cauchy-Dirichlet problem, which is locally uniformly semi-concave in $(0,T)$.
Corollary 1.4.
Let $h$ be a Cauchy-Dirichlet boundary data satisfying (1.3) and (1.4). Then there exists a unique $u\in\mathcal{P}(\Omega_{T})\cap L^{\infty}(\Omega_{T})$ which is locally uniformly semi-concave in $(0,T)$ and satisfy
$\begin{cases}&dt\wedge(dd^{c}u)^{n}=e^{\partial_{t}u+F(t,z,u)}dt\wedge d\mu\text{ in }\Omega_{T},\\ &\lim_{(t,z)\rightarrow(\tau,\zeta)}u(t,z)=h(\tau,\zeta)\text{ for all }(\tau,\zeta)\in[0,T)\times\partial\Omega,\\ &\lim_{t\rightarrow 0+}u(t,\cdot)=h(0,\cdot)\text{ in }L^{1}(d\mu).\end{cases}$
Even though we fixed $0<T<+\infty$, the result of Corollary 1.4 covers the case when $T=+\infty$, which will be explained in Remark 3.10. As an application of Theorem 1.2, we also show the following long-term behavior of the solution under some assumption on $F$.
Theorem 1.5.
Let us denote $T=+\infty$. Let $u\in\mathcal{P}(\Omega_{T})\cap L^{\infty}(\Omega_{T})$ be the unique function which is locally uniformly semi-concave in $(0,T)$ and satisfy
$\begin{cases}&dt\wedge(dd^{c}u)^{n}=e^{\partial_{t}u+F(z,u)}dt\wedge d\mu\text{ in }\Omega_{T},\\ &\lim_{(t,z)\rightarrow(\tau,\zeta)}u(t,z)=0\text{ for all }(\tau,\zeta)\in[0,T)\times\partial\Omega,\\ &\lim_{t\rightarrow 0+}u(t,\cdot)=\psi(\cdot)\text{ in }L^{1}(d\mu),\end{cases}$
where $\psi$ is given in (1.6). If there exists a constant $L_{F}>0$ such that
$\lvert F(z,r_{1})-F(z,r_{2})\rvert\geq L_{F}\lvert r_{1}-r_{2}\rvert$
for all $z\in\Omega$ and $r_{1},r_{2}\in J$, then $u(t,\cdot)\rightarrow\psi_{\infty}$ uniformly as $t\rightarrow+\infty$ for $\psi_{\infty}\in PSH(\Omega)\cap L^{\infty}(\Omega)$ satisfying
$\begin{cases}&(dd^{c}\psi_{\infty})^{n}=e^{F(z,\psi_{\infty})}d\mu\text{ in }\Omega,\\ &\lim_{z\rightarrow\partial\Omega}\psi_{\infty}(z)=0.\end{cases}$
Finally, let us give the historical background of the complex Monge-Ampère flows before we finish the introduction. In 1985, Cao [Cao85] reproved the Yau’s solution to the Calabi conjecture and existence of Kähler-Einstein metric when the first Chern class is negative or zero by using the Kähler-Ricci flow. Since then, the Kähler-Ricci flow has been an important topic in itself with various research subjects, including its behavior on Fano manifolds [PSSW, PS06, Sz10], extension on non-Kähler settings [Gi11, TW15], and so on.
It is a classical result that solving Kähler-Ricci flow is equivalent to solving a parabolic PDE named complex Monge-Ampère flows. After Cao’s result, various parabolic Monge-Ampère equations have been extensively studied, including Monge-Ampère flows, J-flows, inverse Monge-Ampère flows and so on. Recently, Phong and Tô [PT21] introduced the notion of parabolic $C$-subsolutions to study the solvability condition for a large class of parabolic Monge-Ampère equations on compact Hermitian manifolds. Moreover, Picard and Zhang [PZ20] showed the existence and long-time behavior of the solution to the more general parabolic Monge-Ampère equations on compact Kähler manifold without the convexity or concavity condition on the operator, which is extended to the compact Hermitian manifold case by Smith [Sm20].
For the complex Monge-Ampère flows, studying the weak solutions of degenerate equations has been a main interest for its geometrical application, especially for the analytic minimal model program proposed by Song and Tian [ST17]. Eyssidieux, Guedj and Zeriahi [EGZ15, EGZ16] first studied the viscosity approach. Later, Guedj, Lu and Zeriahi [GLZ20, GLZ21-2] introduced the pluripotential approach. They also proved in [GLZ21-1] that both approaches are equivalent under some assumptions. Recently, Dang [Dan24] also extended the pluripotential approach to the Chern-Ricci flow. We refer the reader to [To21, Da22] for the application of viscosity and pluripotential approaches. Besides its geometrical applications, degenerate complex Monge-Ampère flow has been also an interesting topic itself. We refer the reader to [DLT20, DP22] and [Do16, Do17b] for various generalizations of viscosity approach and more general initial data, respectively.
Organization. In Section $2$, we recall the definition of parabolic Monge-Ampère operator in [GLZ21-2]. We also prove several useful lemmas about the operator. In Section $3$, we first prove some lemmas about mixed type inequalities and property of supersolutions. Then we give the proof of Theorem 1.2 in Theorem 3.8 by using these lemmas. With this result, we also prove the uniqueness of the weak solution to the Cauchy-Dirichlet problem and study the long-term behavior of the solution.
Acknowledgements. I would like to thank my advisor, Ngoc Cuong Nguyen, for providing invaluable guidance and support. I would like to thank Quan Tuan Dang for helpful discussion on the proof of the Theorem 3.8 and to Hyunsoo Ahn for a careful reading of the earlier version of its proof. I also thank to Do Hoang Son for useful comments.
2. Parabolic Monge-Ampère Operator
In this section, we prove a comparison principle, a global maximum principle, and a domination principle of the parabolic Monge-Ampère operator. We first recall the definition of the operator in [GLZ21-2].
Definition 2.1.
[GLZ21-2, Definition 2.2] Let $u\in\mathcal{P}(\Omega_{T})\cap L^{\infty}_{loc}(\Omega_{T})$. The map
$C_{c}(\Omega_{T})\ni\chi\mapsto\int_{\Omega_{T}}\chi(t,z)dt\wedge(dd^{c}u)^{n}:=\int_{0}^{T}dt\left(\int_{\Omega}\chi(t,\cdot)(dd^{c}u(t,\cdot))^{n}\right)$
defines a positive Radon measure in $\Omega_{T}$ denoted by $dt\wedge(dd^{c}u)^{n}$.
Note that the integration in the above definition is well-defined since it was proved in [GLZ21-2, Lemma 2.1] that for $u\in\mathcal{P}(\Omega_{T})\cap L^{\infty}_{loc}(\Omega_{T})$ and $\chi\in C_{c}(\Omega_{T})$, the map
$\Gamma_{\chi}:t\mapsto\int_{\Omega}\chi(t,\cdot)(dd^{c}u(t,\cdot))^{n}$
is continuous in $(0,T)$. We first prove a lemma which extends this result.
Lemma 2.2.
Let $u\in\mathcal{P}(\Omega_{T})\cap L^{\infty}_{loc}(\Omega_{T})$ and $E\subset\Omega_{T}$ be a Borel subset. The map
$\Gamma_{E}:t\mapsto\int_{E_{t}}(dd^{c}u(t,\cdot))^{n}$
is Borel measurable in $(0,T)$, where $E_{t}:=\{z\in\Omega~\mid~(t,z)\in E\}$.
Proof.
We first assume that $E$ is compact. Let $\{\chi_{j}(t,z)\}_{j\geq 1}$ be a sequence of continuous positive test functions in $\Omega_{T}$ such that $\chi_{j}\downarrow\mathbf{1}_{E}$ as $j\rightarrow\infty$. Since $\chi_{j}(t,\cdot)\downarrow\mathbf{1}_{E_{t}}$ as $j\rightarrow\infty$ for all $t\in(0,T)$, we have $\Gamma_{\chi_{j}}\downarrow\Gamma_{E}$ as $j\rightarrow\infty$. Therefore $\Gamma_{E}$ is Borel measurable. Next, assume that $E$ is open. Let $\{E_{j}\}_{j\geq 1}$ be a sequence of compact subsets of $\Omega_{T}$ such that $E_{j}\uparrow E$ as $j\rightarrow\infty$. Since $\Gamma_{E_{j}}$ is a Borel measurable function for all $j$ and $\Gamma_{E_{j}}\uparrow\Gamma_{E}$ as $j\rightarrow\infty$, $\Gamma_{E}$ is also Borel measurable.
Finally, we prove the general case. Let us fix open subsets $I\Subset(0,T)$ and $U\Subset\Omega$. Since $u\in\mathcal{P}(\Omega_{T})\cap L^{\infty}_{loc}(\Omega_{T})$, it follows from the elliptic Chern-Levine-Nirenberg inequality (see e.g. [GZ17, Theorem 3.9]) that
(2.1) $\sup_{t\in I}\Gamma_{I\times U}(t)=\sup_{t\in I}\int_{U}(dd^{c}u(t,\cdot))^{n}<+\infty.$
Let $\mathcal{M}$ be a family of all Borel subsets $E$ in $\Omega_{T}$ such that $\Gamma_{E\cap(I\times U)}(t)$ is a Borel measurable function on $(0,T)$. Then $\mathcal{M}$ satisfies the following properties
  • (i)
    $\mathcal{M}$ contains all the open sets in $\Omega_{T}$,
  • (ii)
    If $E,F\in\mathcal{M}$ and $F\subset E$, then $E\setminus F\in\mathcal{M}$,
  • (iii)
    $\mathcal{M}$ is closed under finite disjoint unions,
  • (iv)
    $\mathcal{M}$ is closed under countable increasing unions and countable decreasing intersections.
Since (iii) and (iv) are straightforward, we only explain (i) and (ii). Indeed, (i) is true since if $E\subset\Omega_{T}$ is open, then $E\cap(I\times U)$ is open. Next, if $E,F\in\mathcal{M}$ and $F\subset E$, then
(2.2) $\Gamma_{(E\setminus F)\cap(I\times U)}(t)$ $=\int_{(E\setminus F)_{t}\cap U}(dd^{c}u(t,\cdot))^{n}$
$=\int_{E_{t}\cap U}(dd^{c}u(t,\cdot))^{n}-\int_{F_{t}\cap U}(dd^{c}u(t,\cdot))^{n}$
$=\Gamma_{E\cap(I\times U)}(t)-\Gamma_{F\cap(I\times U)}(t).$
It implies that $E\setminus F\in\mathcal{M}$. Note that we used (2.1) for the second equality of (2.2). Therefore by proceeding as the proof of [Fo84, Theorem 7.26], one can prove that any Borel subset $E$ of $\Omega_{T}$ is contained in $\mathcal{M}$. Now, let $\{I_{j}\}_{j\geq 1}$ and $\{U_{j}\}_{h\geq 1}$ be sequence of relatively compact open subsets of $(0,T)$ and $\Omega$ such that $I_{j}\uparrow(0,T)$ and $U_{j}\uparrow\Omega$. If $E$ is a Borel subset of $\Omega_{T}$, then by the preceding arguments, $\Gamma_{E\cap(I_{j}\times U_{j})}(t)$ is a Borel measurable function on $(0,T)$ for all $j$. By taking $j\rightarrow\infty$, we get the conclusion. ∎
From now, we use the notation $E_{t}$ for a Borel set $E\subset\Omega_{T}$ as defined in Lemma 2.2 throughout this paper. Note that because of Lemma 2.2, the integration
(2.3) $\int_{0}^{T}dt\int_{E_{t}}(dd^{c}u(t,\cdot))^{n}$
is well-defined for all Borel subsets $E\subset\Omega_{T}$. We next show that (2.3) coincides with the integration of $E$ with respect to a Radon measure $dt\wedge(dd^{c}u)^{n}$, which is defined in Definition 2.1. We will use this lemma to compute and compare the size of sublevel sets with respect to the measures of the form $dt\wedge(dd^{c}u)^{n}$ where $u\in\mathcal{P}(\Omega_{T})\cap L^{\infty}(\Omega_{T})$.
Lemma 2.3.
Let $u\in\mathcal{P}(\Omega_{T})\cap L^{\infty}_{loc}(\Omega_{T})$. For any Borel subset $E\subset\Omega_{T}$, we have
$\int_{E}dt\wedge(dd^{c}u)^{n}=\int_{0}^{T}dt\int_{E_{t}}(dd^{c}u(t,\cdot))^{n}.$
Proof.
Let us first assume that $E$ is compact. Let $J\subset(0,T)$ and $K\subset\Omega$ be compact subsets such that $E\subset J\times K$. It follows from the elliptic Chern-Levine-Nirenberg inequality that
(2.4) $\sup_{t\in(0,T)}\int_{E_{t}}(dd^{c}u(t,\cdot))^{n}\leq\sup_{t\in J}\int_{K}(dd^{c}u(t,\cdot))^{n}<+\infty.$
Let $\{\chi_{j}(t,z)\}_{j\geq 1}$ be a sequence of continuous positive test functions in $\Omega_{T}$ such that $supp(\chi_{j})\subset J\times K$ for all $j$ and $\chi_{j}\downarrow\mathbf{1}_{E}$ as $j\rightarrow\infty$. By the definition of $dt\wedge(dd^{c}u)^{n}$,
(2.5) $\int_{\Omega_{T}}\chi_{j}(t,z)dt\wedge(dd^{c}u)^{n}=\int_{0}^{T}dt\int_{\Omega}\chi_{j}(t,\cdot)(dd^{c}u(t,\cdot))^{n}.$
Since $\chi_{j}(t,\cdot)\downarrow\mathbf{1}_{E_{t}}$ as $j\rightarrow\infty$ for all $t\in(0,T)$, it follows from (2.4) and the dominated convergence theorem that
(2.6) $\lim_{j\rightarrow\infty}\int_{0}^{T}dt\int_{\Omega}\chi_{j}(t,\cdot)(dd^{c}u(t,\cdot))^{n}=\int_{0}^{T}dt\int_{E_{t}}(dd^{c}u(t,\cdot))^{n}.$
Combining (2.5) and (2.6), we have
$\int_{E}dt\wedge(dd^{c}u)^{n}=\lim_{j\rightarrow\infty}\int_{\Omega_{T}}\chi_{j}(t,z)dt\wedge(dd^{c}u)^{n}=\int_{0}^{T}dt\int_{E_{t}}(dd^{c}u(t,\cdot))^{n}.$
Next, for an open subset $E\subset\Omega_{T}$, let $\{E_{j}\}_{j\geq 1}$ be a sequence of compact subsets of $\Omega_{T}$ such that $E_{j}\uparrow E$ as $j\rightarrow\infty$. It follows from the monotone convergence theorem that
$\lim_{j\rightarrow\infty}\int_{E_{j}}dt\wedge(dd^{c}u)^{n}=\int_{E}dt\wedge(dd^{c}u)^{n}$
and
$\lim_{j\rightarrow\infty}\int_{0}^{T}dt\int_{(E_{j})_{t}}(dd^{c}u(t,\cdot))^{n}=\int_{0}^{T}dt\int_{E_{t}}(dd^{c}u(t,\cdot))^{n}.$
Therefore we get the result since we have
$\int_{E_{j}}dt\wedge(dd^{c}u)^{n}=\int_{0}^{T}dt\int_{(E_{j})_{t}}(dd^{c}u(t,\cdot))^{n}$
for all $j$. Finally, for a Borel subset $E\subset\Omega_{T}$, one can prove the result by using inner and outer regularity of the Radon measure $dt\wedge(dd^{c}u)^{n}$. ∎
We now prove a parabolic analogue of the classical comparison principle.
Lemma 2.4.
Let $u,v\in\mathcal{P}(\Omega_{T})\cap L^{\infty}(\Omega_{T})$ satisfying $\liminf_{(t,z)\rightarrow[0,T)\times\partial\Omega}(v(t,z)-u(t,z))\geq 0$. Then
$\int_{\{v<u\}}dt\wedge(dd^{c}u)^{n}\leq\int_{\{v<u\}}dt\wedge(dd^{c}v)^{n}.$
Proof.
Fix $\varepsilon_{1}>0$ and $0<T^{\prime}<T$. Since $\liminf_{(t,z)\rightarrow[0,T^{\prime}]\times\partial\Omega}(v(t,z)-u(t,z))\geq 0$, there exists a compact set $K\subset\Omega$ such that
$\{v<u-\varepsilon_{1}\}\cap((0,T^{\prime})\times\Omega)\subset[0,T^{\prime}]\times K.$
Let $\chi_{\varepsilon_{2}}(z)$ be a standard smooth mollifier and define
$u_{\varepsilon_{2}}(t,z):=u(t,z)*\chi_{\varepsilon_{2}}(z)=\int_{\Omega}u(t,\zeta)\chi_{\varepsilon_{2}}(z-\zeta)dV(\zeta).$
For sufficiently small $\varepsilon_{2}>0$, $u_{\varepsilon_{2}}$ is defined on $(0,T)\times\Omega^{\prime}$ for some open set $\Omega^{\prime}$ satisfying $K\Subset\Omega^{\prime}\Subset\Omega$. Moreover, one can check that $u_{\varepsilon_{2}}\in\mathcal{P}((0,T)\times\Omega^{\prime})\cap C((0,T)\times\Omega^{\prime})$ and $u_{\varepsilon_{2}}\downarrow u$ as $\varepsilon_{2}\rightarrow 0$. Since $v$ is (jointly) upper semi-continuous by [GLZ21-2, Proposition 1.4], $\{v<u_{\varepsilon_{2}}-\varepsilon_{1}\}\cap((0,T^{\prime})\times\Omega^{\prime})$ is an open subset of $\Omega_{T}$ which increases to $\{v<u-\varepsilon_{1}\}\cap((0,T^{\prime})\times\Omega)$ as $\varepsilon_{2}\rightarrow 0$. By taking $\varepsilon_{1}\rightarrow 0$ and $T^{\prime}\rightarrow T$, one can see that $\{v<u\}$ is a Borel subset of $\Omega_{T}$.
It follows from the classical comparison principle (see e.g. [GZ17, Theorem 3.29]) that for all $t\in(0,T)$,
(2.7) $\int_{\{v(t,\cdot)<u(t,\cdot)\}}(dd^{c}u(t,\cdot))^{n}\leq\int_{\{v(t,\cdot)<u(t,\cdot)\}}(dd^{c}v(t,\cdot))^{n}.$
It follows from Lemma 2.2 that each side of (2.7) is a Borel measurable function on $(0,T)$. By integrating each side with respect to $dt$ for $t\in(0,T)$,
$\int_{0}^{T}dt\int_{\{v(t,\cdot)<u(t,\cdot)\}}(dd^{c}u(t,\cdot))^{n}\leq\int_{0}^{T}dt\int_{\{v(t,\cdot)<u(t,\cdot)\}}(dd^{c}v(t,\cdot))^{n}.$
It follows from Lemma 2.3 that
$\int_{\{v<u\}}dt\wedge(dd^{c}u)^{n}\leq\int_{\{v<u\}}dt\wedge(dd^{c}v)^{n},$
which completes the proof. ∎
The following two lemmas are parabolic analogues of classical global maximum principle and domination principle. The proofs are almost same with the ones in [GZ17, Corollary 3.30, Corollary 3.31].
Lemma 2.5.
Let $u,v\in\mathcal{P}(\Omega_{T})\cap L^{\infty}(\Omega_{T})$ satisfying $\liminf_{(t,z)\rightarrow[0,T)\times\partial\Omega}(v(t,z)-u(t,z))\geq 0$. If $dt\wedge(dd^{c}v)^{n}\leq dt\wedge(dd^{c}u)^{n}$, then $u\leq v$ in $\Omega_{T}$.
Proof.
For $\varepsilon>0$, we set $\phi_{\varepsilon}:=u+\varepsilon\rho$, where $\rho(t,z):=\lvert z\rvert^{2}-R^{2}$ for $R>0$ large enough so that $\rho<0$ on $\Omega_{T}$. Then $\phi_{\varepsilon}\in\mathcal{P}(\Omega_{T})\cap L^{\infty}(\Omega_{T})$ and $\liminf_{z\rightarrow\partial\Omega}(v(t,z)-\phi_{\varepsilon}(t,z))\geq 0$ for all $t\in(0,T)$. It follows from Lemma 2.4 that
$\int_{\{v<\phi_{\varepsilon}\}}dt\wedge(dd^{c}{\phi}_{\varepsilon})^{n}\leq\int_{\{v<\phi_{\varepsilon}\}}dt\wedge(dd^{c}v)^{n}.$
Since
$dt\wedge(dd^{c}\phi_{\varepsilon})^{n}\geq dt\wedge(dd^{c}v)^{n}+\varepsilon^{n}dt\wedge(dd^{c}\rho)^{n},$
we have $\int_{\{v<\phi_{\varepsilon}\}}dt\wedge(dd^{c}\rho)^{n}=0$. It implies that the set $\{v<\phi_{\varepsilon}\}$ has Lebesgue measure zero, so that $\{v<u\}$ also has Lebesgue measure 0. Therefore by [GLZ21-2, Corollary 1.9], we have $u\leq v$ in $\Omega_{T}$. ∎
Lemma 2.6.
Let $u,v\in\mathcal{P}(\Omega_{T})\cap L^{\infty}(\Omega_{T})$ satisfying $\liminf_{(t,z)\rightarrow[0,T)\times\partial\Omega}(v(t,z)-u(t,z))\geq 0$. Assume that for all $u\leq v$ a.e. in $\Omega_{T}$ with respect to the measure $dt\wedge(dd^{c}v)^{n}$. Then $u\leq v$ in $\Omega_{T}$.
Proof.
For $\varepsilon>0$, we set $u_{\varepsilon}:=\phi_{\varepsilon}+\varepsilon\rho$, where $\rho(t,z):=\lvert z\rvert^{2}-R^{2}$ for $R>0$ large enough so that $\rho<0$ on $\Omega_{T}$. By Lemma 2.4,
$\int_{\{v<\phi_{\varepsilon}\}}dt\wedge(dd^{c}\phi_{\varepsilon})^{n}\leq\int_{\{v<\phi_{\varepsilon}\}}dt\wedge(dd^{c}v)^{n}\leq\int_{\{v<u\}}dt\wedge(dd^{c}v)^{n}=0.$
Since $dt\wedge(dd^{c}\phi_{\varepsilon})^{n}\geq\varepsilon^{n}dt\wedge(dd^{c}\rho)^{n}$, it implies that the set $\{v<\phi_{\varepsilon}\}$ has Lebesgue measure 0, so that $\{v<u\}$ also has Lebesgue measure 0. Therefore by [GLZ21-2, Corollary 1.9], we have $u\leq v$ in $\Omega_{T}$. ∎
3. A Generalized Comparison Principle
In this section, we give a proof of Theorem 1.2. We first prove some useful lemmas which will be used to prove it.
3.1. Mixed type inequalities
We show the following lemmas on mixed type inequalities. The main result of [GLZ19] was used to characterize a subsolution in the last part of the proof of [GLZ21-2, Theorem 6.6], which is not available in our case. The following lemmas will be used instead.
Lemma 3.1.
Assume that $\varphi,\psi\in PSH(\Omega)\cap L^{\infty}(\Omega)$ satisfy
$(dd^{c}\varphi)^{n}\geq f(dd^{c}\psi)^{n}$
for some $f\in L^{1}(\Omega,(dd^{c}\psi)^{n})$. Then, we have
$(dd^{c}\varphi)\wedge(dd^{c}\psi)^{n-1}\geq f^{1/n}(dd^{c}\psi)^{n}.$
Proof.
It follows directly from Dinew’s mixed type inequality in [Di09, Theorem 1.3]. ∎
Lemma 3.2.
Assume that $u,\varphi\in PSH(\Omega)\cap L^{\infty}(\Omega)$ satisfy $\liminf_{z\rightarrow\partial\Omega}(\varphi(z)-u(z))\geq 0$. Then
$\int_{\{\varphi<u\}}(dd^{c}u)\wedge(dd^{c}\varphi)^{n-1}\leq\int_{\{\varphi<u\}}(dd^{c}\varphi)^{n}.$
Proof.
The proof is almost same with the one for the comparison principle by Bedford and Taylor in [BT82, Theorem 4.2] (see also [Ko05, Theorem 1.16]). ∎
3.2. Boundary behavior of supersolutions
Next, we show the following lemma about the property of supersolutions. Let $h$ be a Cauchy-Dirichlet boundary data and denote $v_{0}(z):=h(0,z)$ for $z\in\Omega$.
Lemma 3.3.
Let us assume that $v\in\mathcal{P}(\Omega_{T})\cap L^{\infty}(\Omega_{T})$ satisfy
$\begin{cases}&dt\wedge(dd^{c}v)^{n}\leq e^{\partial_{t}v+F(t,z,v)}dt\wedge d\mu,\\ &\lim_{(t,z)\rightarrow(\tau,\zeta)\in[0,T)\times\partial\Omega}v(t,z)=h(\tau,\zeta),\\ &\lim_{t\rightarrow 0+}v(t,\cdot)=v_{0}\text{ in }L^{1}(d\mu).\end{cases}$
Then there exists a function $\eta(t)\in C([0,T))$ such that $\eta(0)=0$ and
$v(t,z)\geq v_{0}(z)-\eta(t)$
for $dt\wedge d\mu$-a.e. $(t,z)\in\Omega_{T}$.
Proof.
We proceed as the proof of [GLZ21-2, Lemma 6.9]. Let us fix $0<S<T$. For $s>0$ small enough, we set
$\delta(s):=\sup\{\lvert h(\tau,z)-h(t,z)\rvert~\mid~z\in\partial\Omega,~t,\tau\in[0,S],~\lvert t-\tau\rvert\leq s\}.$
Since $h$ is continuous on $0,T)\times\partial\Omega$, $\lim_{s\rightarrow 0+}\delta(s)=0$.
Now, fix $0<s<\frac{S}{2}$. We show that for some constant $C>0$,
$v(s,z)\geq v_{0}(z)-\delta(s)+s(\psi(z)-C)+n(s\log(s/T)-s)$
holds for $\mu$-a.e. $z\in\Omega$. Note that $\psi$ is given in (1.6).
Fix $0<\varepsilon\leq s$. Let us define a map $(t,z)\mapsto U^{\varepsilon}(t,z)$ on $\Omega_{s}$ by
$U^{\varepsilon}(t,z):=v(\varepsilon,z)-\delta(s)+t(\psi(z)-C_{1})+n(t\log(t/T)-t)$
for some constant $C_{1}>0$ so that $U^{\varepsilon}$ satisfies
$dt\wedge(dd^{c}U^{\varepsilon})^{n}\geq e^{\partial_{t}U^{\varepsilon}+F(t,z,U^{\varepsilon})}dt\wedge d\mu\text{ in }\Omega_{s}.$
Next, let us define a map $(t,z)\mapsto V^{\varepsilon}(t,z)$ on $\Omega_{s}$ by
$V^{\varepsilon}(t,z):=v(t+\varepsilon,z)+C_{2}\varepsilon t$
for some constant $C_{2}>0$ so that $V^{\varepsilon}$ satisfies
$dt\wedge(dd^{c}V^{\varepsilon})^{n}\leq e^{\partial_{t}V^{\varepsilon}+F(t,z,V^{\varepsilon})}dt\wedge d\mu\text{ in }\Omega_{s}.$
Note that $U^{\varepsilon}$ and $V^{\varepsilon}$ satisfy the following properties
  • (i)
    $\liminf_{z\rightarrow\partial\Omega}(V^{\varepsilon}(t,z)-U^{\varepsilon}(t,z))\geq 0$ for all $t\in(0,T)$,
  • (ii)
    $\lim_{t\rightarrow 0+}U^{\varepsilon}(t,z)=v(\varepsilon,z)-\delta(s)$ uniformly for all $z\in\Omega$,
  • (iii)
    $\lim_{t\rightarrow 0+}V^{\varepsilon}(t,z)=v(\varepsilon,z)$ uniformly for all $z\in\Omega$.
Also, $U^{\varepsilon}$ and $V^{\varepsilon}$ are locally uniformly semi-concave in $(0,s)$, and
$\partial_{t}U^{\varepsilon}(t,z)=\psi(z)-C_{1}+n\log(t/T),$
which is locally uniformly Lipschitz in $(0,T)$. Therefore it follows from Lemma 3.4 that $V^{\varepsilon}(t,z)\geq U^{\varepsilon}(t,z)$ for all $(t,z)\in\Omega_{s}$. Recall that $\lim_{t\rightarrow 0+}v(t,\cdot)=v_{0}$ in $L^{1}(d\mu)$. Hence there exists a sequence $\{\varepsilon_{k}\}_{k\geq 1}$ such that $\varepsilon_{k}\downarrow 0$ as $k\rightarrow\infty$ and $v(t_{k},z)\rightarrow v(0,z)$ as $k\rightarrow\infty$ for $d\mu$-a.e. $z\in\Omega$. Hence after substituting $\varepsilon$ by $\varepsilon_{k}$ and taking $k\rightarrow\infty$, we get
$v(t,z)\geq v_{0}(z)-\delta(s)+t(\psi(z)-C_{1})+n(t\log(t/T)-t)$
holds for all $t\in(0,s)$ and $\mu$-a.e. $z\in\Omega$. By taking $t\rightarrow s$, we get the result. ∎
3.3. Main result
We now introduce a lemma which plays a key role in this paper. Recall that for a bounded parabolic potential $u$ which is locally uniformly semi-concave in $(0,T)$,
$\partial_{t}^{+}u(t,z):=\lim_{\delta\rightarrow 0+}\frac{u(t+\delta,z)-u(t,z)}{\delta}$
and
$\partial_{t}^{-}u(t,z):=\lim_{\delta\rightarrow 0-}\frac{u(t+\delta,z)-u(t,z)}{\delta}$
exist for all $(t,z)\in\Omega_{T}$ and $\partial_{t}u(t,z)$ exists when $\partial_{t}^{+}u(t,z)=\partial_{t}^{-}u(t,z)$.
Lemma 3.4.
Assume that $u,v\in\mathcal{P}(\Omega_{T})\cap L^{\infty}(\Omega_{T})$ satisfy
  1. (a)
    $\liminf_{(t,z)\rightarrow[0,T)\times\partial\Omega}(v(t,z)-u(t,z))\geq 0$,
  2. (b)
    for each $\varepsilon>0$, there exists $\tilde{\varepsilon}>0$ such that
    $\int_{\{v+(t+1)\varepsilon<u\}\cap((0,\tilde{\varepsilon})\times\Omega)}dt\wedge d\mu=0,$
  3. (c)
    $u$ and $v$ are locally uniformly semi-concave in $(0,T)$,
  4. (d)
    for a.e. $t\in(0,T)$,
    $(dd^{c}u(t,\cdot))\wedge(dd^{c}v(t,\cdot))^{n-1}$
    $\quad\geq\exp\left(\frac{\partial_{t}u(t,\cdot)-\partial_{t}v(t,\cdot)+F(t,\cdot,u)-F(t,\cdot,v)}{n}\right)(dd^{c}v(t,\cdot))^{n},$
  5. (e)
    $dt\wedge(dd^{c}v)^{n}\leq e^{\partial_{t}v+F(t,z,v)}dt\wedge d\mu$ in $\Omega_{T}$,
  6. (f)
    $\partial_{t}u(t,z)$ is well-defined for all $(t,z)\in\Omega_{T}$ and locally uniformly Lipschitz in $(0,T)$.
Then $u\leq v$ in $\Omega_{T}$.
Proof.
Fix $\varepsilon>0$. Let us define
$E:=\{v+(t+1)\varepsilon<u\}.$
Our goal is to show that $E=\emptyset$. Assume that it is not the case. Let us denote $E_{t}:=\{z\in\Omega~\mid~(t,z)\in E\}$. We define
$t_{1}:=\inf\{t\in(0,T)~\mid~\mu(E_{t})>0\}.$
The following Figure 1 describes our strategy of the proof.
Figure 1.
The idea is to prove that $t_{1}>0$ as in the left figure, and then compute the size of the set $E\cap((t_{1}-\delta,t_{1}+\delta)\times\Omega)$ which is described in the right figure. We will show that it is actually an empty set, which contradicts to the definition of $t_{1}$.
We first show that $t_{1}>0$. By the assumptions (b) and (e), there exists $\tilde{\varepsilon}>0$ such that
$0\leq\int_{E\cap((0,\tilde{\varepsilon})\times\Omega)}dt\wedge(dd^{c}v)^{n}\leq\int_{E\cap((0,\tilde{\varepsilon})\times\Omega)}e^{\partial_{t}v+F(t,z,v)}dt\wedge d\mu=0,$
which implies that
$\int_{E\cap((0,\tilde{\varepsilon})\times\Omega)}dt\wedge(dd^{c}v)^{n}=0.$
[…]
Next, by the assumption (f), there exists $C_{1}>0$ such that
$\lvert\partial_{t}u(t,z)-\partial_{t}u(s,z)\rvert<C_{1}\lvert t-s\rvert$
for all $t,s\in[t_{1}-\tilde{\varepsilon},t_{1}+\tilde{\varepsilon}$] and $z\in\Omega$. It follows from the assumption (c) that there exists $C_{2}>0$ such that $u(t,z)-C_{2}t^{2}$ and $v(t,z)-C_{2}t^{2}$ is concave in $t_{1}-\tilde{\varepsilon},t_{1}+\tilde{\varepsilon}$] for all $z\in\Omega$. Let us define $C_{3}:=\max\{C_{1},C_{2}\}$, so that we have
(3.1) $\lvert\partial_{t}u(t,z)-\partial_{t}u(s,z)\rvert<C_{3}\lvert t-s\rvert$
for all $t,s\in[t_{1}-\tilde{\varepsilon},t_{1}+\tilde{\varepsilon}$] and $z\in\Omega$, and
(3.2) $u(t,z)-C_{3}t^{2},v(t,z)-C_{3}t^{2}\text{ are concave in }[t_{1}-\tilde{\varepsilon},t_{1}+\tilde{\varepsilon}$]
for all $z\in\Omega.$
Let us denote $\delta:=\min\{\frac{\tilde{\varepsilon}}{2},\frac{\varepsilon}{6C_{3}}\}$ and fix $t\in(t_{1}-\delta,t_{1}+\delta)$. Since $\mu(E_{t-\delta})=0$, we have
$u(t,z)>v(t,z)+(t+1)\varepsilon\text{ for all }z\in E_{t}$
and
$u(t-\delta,z)\leq v(t-\delta,z)+(t-\delta+1)\varepsilon\text{ for $\mu$-a.e. }z\in E_{t}.$
Therefore we have
(3.3) $\frac{u(t,z)-u(t-\delta,z)}{\delta}>\frac{v(t,z)-v(t-\delta,z)}{\delta}+\varepsilon\text{ for $\mu$-a.e. $z\in E_{t}$.}$
It follows from (3.2) that
$\frac{(u(t,z)-C_{3}t^{2})-(u(t-\delta,z)-C_{3}(t-\delta)^{2})}{\delta}$ $\leq\partial_{t}^{-}u(t-\delta,z)-2C_{3}(t-\delta)$
$\quad=\partial_{t}u(t-\delta,z)-2C_{3}(t-\delta)$
and
$\frac{(v(t,z)-C_{3}t^{2})-(v(t-\delta,z)-C_{3}(t-\delta)^{2})}{\delta}\geq\partial_{t}^{-}v(t,z)-2C_{3}t$
for all $z\in E_{t}$. Hence one may induce from (3.3) that
(3.4) $\partial_{t}u(t-\delta,z)+2C_{3}\delta>\partial_{t}^{-}v(t,z)+\varepsilon\text{ for $\mu$-a.e. $z\in E_{t}$.}$
[…]
(3.5) $\partial_{t}u(t,z)>\partial_{t}^{-}v(t,z)+\frac{\varepsilon}{2}\text{ for $\mu$-a.e. $z\in E_{t}$.}$
We now show that $E\cap((t_{1}-\delta,t_{1}+\delta)\times\Omega)=\emptyset$. For a.e. $t\in(t_{1}-\delta,t_{1}+\delta)$, we have
(3.6) $\int_{E(t,\cdot)}e^{\varepsilon/2n}(dd^{c}v(t,\cdot))^{n}$
$\quad\leq\int_{E(t,\cdot)}\exp\left(\frac{\partial_{t}u(t,\cdot)-\partial_{t}v(t,\cdot)+F(t,\cdot,u)-F(t,\cdot,v)}{n}\right)(dd^{c}v(t,\cdot))^{n}$
$\quad\leq\int_{E(t,\cdot)}(dd^{c}u(t,\cdot))\wedge(dd^{c}v(t,\cdot))^{n-1}$
$\quad\leq\int_{E(t,\cdot)}(dd^{c}v(t,\cdot))^{n},$
where we used (3.5) for the first inequality and Lemma 3.2 for the last inequality. Indeed, we could use (3.5) since
$(dd^{c}v(t,\cdot))^{n}\leq e^{\partial_{t}v(t,\cdot)+F(t,\cdot,v)}d\mu$
for a.e. $t\in(0,T)$ by [GLZ21-2, Proposition 3.2], which implies that for a.e. $t\in(t_{1}-\delta,t_{1}+\delta)$, (3.5) holds for $(dd^{c}v(t,\cdot))^{n}$-a.e. $z\in E_{t}$. By integrating both sides of (3.6) with respect to $t$ for $(t_{1}-\delta,t_{1}+\delta)$, we get
$0\leq e^{\varepsilon/2n}\int_{t_{1}-\delta}^{t_{1}+\delta}dt\int_{E(t,\cdot)}(dd^{c}v(t,\cdot))^{n}\leq\int_{t_{1}-\delta}^{t_{1}+\delta}dt\int_{E(t,\cdot)}(dd^{c}v(t,\cdot))^{n},$
which implies that
$\int_{t_{1}-\delta}^{t_{1}+\delta}dt\int_{E(t,\cdot)}(dd^{c}v(t,\cdot))^{n}=0.$
It follows from Lemma 2.3 that
$\int_{E\cap((t_{1}-\delta,t_{1}+\delta)\times\Omega)}dt\wedge(dd^{c}v)^{n}=\int_{t_{1}-\delta}^{t_{1}+\delta}dt\int_{E(t,\cdot)}(dd^{c}v(t,\cdot))^{n}=0.$
This implies that $E\cap((t_{1}-\delta,t_{1}+\delta)\times\Omega)=\emptyset$ by Lemma 2.6. Since it contradicts to the definition of $t_{1}$, we conclude that $E=\emptyset$. ∎
Remark 3.5.
We provide an example for which the assumptions of the above lemma are satisfied. First, $u$ and $v$ satisfy the condition (b) if $\lim_{t\rightarrow 0+}v(t,z)=v_{0}(z)$ and $\lim_{t\rightarrow 0+}u(t,z)=u_{0}(z)$ uniformly where $v_{0},u_{0}\in PSH(\Omega)\cap L^{\infty}(\Omega)$ satisfy $v_{0}\geq u_{0}$. Indeed, for all $z\in\Omega$, there exists $\delta_{1}>0$ such that if $0<\lvert t\rvert<\delta_{1}$, then $\lvert v_{0}(z)-v(t,z)\rvert<\varepsilon/2$ and $\lvert u_{0}(z)-u(t,z)\rvert<\varepsilon/2$. Therefore for all $(t,z)\in(0,\delta_{1})\times\Omega$,
$v(t,z)+(t+1)\varepsilon$ $>v_{0}(z)+\left(t+\frac{1}{2}\right)\varepsilon$
$\geq u_{0}(z)+\left(t+\frac{1}{2}\right)\varepsilon$
$>u(t,z)+t\varepsilon$
$>u(t,z).$
Next, $u$ and $v$ satisfy the condition (d) if $dt\wedge(dd^{c}u)^{n}\geq e^{\partial_{t}u+F(t,z,u)}dt\wedge d\mu$ and $dt\wedge(dd^{c}v)^{n}\leq e^{\partial_{t}v+F(t,z,v)}dt\wedge d\mu$. Indeed, by using [GLZ21-2, Proposition 3.2], one can show that for a.e. $t\in(0,T)$, $u$ and $v$ satisfy
$(dd^{c}u(t,\cdot))^{n}\geq e^{\partial_{t}u(t,\cdot)+F(t,\cdot,u)}d\mu\geq e^{\partial_{t}u(t,\cdot)-\partial_{t}v(t,\cdot)+F(t,\cdot,u)-F(t,\cdot,v)}(dd^{c}v(t,\cdot))^{n}.$
It follows from Lemma 3.1 that for a.e. $t\in(0,T)$,
$(dd^{c}u(t,\cdot))\wedge(dd^{c}v(t,\cdot))^{n-1}$
$\quad\geq\exp\left(\frac{\partial_{t}u(t,\cdot)-\partial_{t}v(t,\cdot)+F(t,\cdot,u)-F(t,\cdot,v)}{n}\right)(dd^{c}v(t,\cdot))^{n}.$
We next prove a lemma about the condition (b) on Lemma 3.4. This will be used later while proving Theorem 1.2.
Lemma 3.6.
Assume that $u,v\in\mathcal{P}(\Omega_{T})\cap L^{\infty}(\Omega_{T})$ satisfy
  1. (a)
    $\liminf_{(t,z)\rightarrow[0,T)\times\partial\Omega}(v(t,z)-u(t,z))\geq 0$,
  2. (b)
    there exists a function $\eta(t)\in C([0,T])$ such that $\eta(0)=0$ and
    $v(t,z)\geq v_{0}(z)-\eta(t)$
    for $dt\wedge d\mu$-a.e. $(t,z)\in\Omega_{T}$,
  3. (c)
    for each open $U\Subset\Omega$, there exists $A>0$ such that
    $\int_{U}u(t,\cdot)d\mu\leq\int_{U}u_{0}d\mu+At,$
  4. (d)
    $u_{0}(z)\leq v_{0}(z)$ for all $z\in\Omega$.
Then for each $\varepsilon>0$, there exists $\tilde{\varepsilon}>0$ such that
$\int_{\{v+(t+1)\varepsilon<u\}\cap((0,\tilde{\varepsilon})\times\Omega)}dt\wedge d\mu=0.$
Proof.
We fix $\varepsilon>0$. Let us define
$E:=\{v+(t+1)\varepsilon<u\}.$
Let us fix $0<S<T$. Let $U\Subset\Omega$ be an open subset satisfying $E\cap((0,S]\times\Omega)\subset[0,S]\times U$. Since
$1<\frac{u-v}{(t+1)\varepsilon}$
in $E$, for all $0<\delta<S$, we have
(3.7) $\int_{E\cap((0,\delta)\times U)}dt\wedge d\mu$
$\quad\leq\int_{E\cap((0,\delta)\times U)}\frac{u-v}{(t+1)\varepsilon}dt\wedge d\mu$
$\quad\leq\int_{E\cap((0,\delta)\times U)}\frac{u-u_{0}}{(t+1)\varepsilon}dt\wedge d\mu+\int_{E\cap((0,\delta)\times U)}\frac{v_{0}-v}{(t+1)\varepsilon}dt\wedge d\mu$
$\quad\leq\frac{A}{\varepsilon}\int_{E\cap((0,\delta)\times U)}\frac{t}{t+1}dt\wedge d\mu+\frac{1}{\varepsilon}\int_{E\cap((0,\delta)\times U)}\frac{\eta(t)}{t+1}dt\wedge d\mu$
$\quad\leq\frac{A\delta+\sup_{t\in(0,\delta)}\lvert\eta(t)\rvert}{\varepsilon}\int_{E\cap((0,\delta)\times U)}dt\wedge d\mu.$
Here, we used the condition (d) for the second inequality and the condition (b) and (c) for the third inequality. Let us choose $\tilde{\varepsilon}$ sufficiently small so that
(3.8) $A\tilde{\varepsilon}\leq\frac{\varepsilon}{3}\text{ and }\sup_{t\in(0,\tilde{\varepsilon})}\lvert\eta(t)\rvert\leq\frac{\varepsilon}{3}.$
By combining (3.7) and (3.8), we get
$0\leq\int_{E\cap((0,\tilde{\varepsilon})\times U)}dt\wedge d\mu\leq\frac{2}{3}\int_{E\cap((0,\tilde{\varepsilon})\times U)}dt\wedge d\mu,$
which implies that
$\int_{E\cap((0,\tilde{\varepsilon})\times\Omega)}dt\wedge d\mu=\int_{E\cap((0,\tilde{\varepsilon})\times U)}dt\wedge d\mu=0.$
Thus we get the result. ∎
Remark 3.7.
We do not need (1.5) for the Lemma 3.4 and 3.6. They are indeed true for any positive Borel measure $\mu$ satisfying $PSH(\Omega)\subset L^{1}_{loc}(\Omega,d\mu)$.
Finally, we are ready to prove our main result.
Theorem 3.8.
Let $u,v\in\mathcal{P}(\Omega_{T})\cap L^{\infty}(\Omega_{T})$ with the Cauchy-Dirichlet boundary data $h_{1}$, $h_{2}$ as in (1.2). Assume that
  1. (a)
    $u$ and $v$ are locally uniformly semi-concave in $(0,T)$;
  2. (b)
    $dt\wedge(dd^{c}u)^{n}\geq e^{\partial_{t}u+F(t,z,u)}dt\wedge d\mu$ in $\Omega_{T}$;
  3. (c)
    $dt\wedge(dd^{c}v)^{n}\leq e^{\partial_{t}v+F(t,z,v)}dt\wedge d\mu$ in $\Omega_{T}$;
  4. (d)
    $h_{1}$ satisfies (1.3).
If $h_{1}\leq h_{2}$, then $u\leq v$.
Proof.
Fix $0<S<T$. We first assume that
(3.9) $t\lvert\partial_{t}u(t,z)\rvert\leq B$
for some $B>0$ in $\Omega_{S}$. We will remove this assumption at the end of the proof. For $s>0$ near $1$ we set, for $(t,z)\in\Omega_{S}$,
$W^{s}(t,z):=s^{-1}u(st,z)-C\lvert s-1\rvert(t+1).$
It follows from [GLZ21-2, Theorem 4.2] that for $C>0$ large enough, we have
(3.10) $\begin{cases}&dt\wedge(dd^{c}W^{s})^{n}\geq e^{\partial_{t}W^{s}+F(t,z,W^{s})}dt\wedge d\mu\text{ in }\Omega_{S},\\ &\limsup_{(t,z)\rightarrow(\tau,\zeta)\in[0,S)\times\partial\Omega}W^{s}(t,z)\leq h_{1}(\tau,\zeta),\\ &\limsup_{t\rightarrow 0+}W^{s}(t,z)\leq h_{1}(0,z)\text{ for $\mu$-a.e. $z\in\Omega$.}\end{cases}$
Note that we need the condition (d) to find such a constant $C>0$. Let $\eta_{\varepsilon}$ be the standard mollifier on $\mathbb{R}$. We set
$\Phi^{\varepsilon}(t,z):=\int_{\mathbb{R}}W^{s}(t,z)\eta_{\varepsilon}(s-1)ds.$
The idea is to apply the Lemma 3.4 for $\Phi^{\varepsilon}-O(\varepsilon)t$ and $v$ in $\Omega_{S}$, and derive the result. We show that they satisfy all the assumptions given in Lemma 3.4.
One can check that $\partial_{t}\Phi^{\varepsilon}(t,z)$ is well-defined for all $(t,z)\in\Omega$ and $\Phi^{\varepsilon}$ satisfies
$\begin{cases}&\limsup_{(t,z)\rightarrow(\tau,\zeta)\in[0,S)\times\partial\Omega}\Phi^{\varepsilon}(t,z)\leq h_{1}(\tau,\zeta),\\ &\limsup_{t\rightarrow 0+}\Phi^{\varepsilon}(t,z)\leq h_{1}(0,z)\text{ for $\mu$-a.e. $z\in\Omega$.}\end{cases}$
We prove that $\partial_{t}\Phi^{\varepsilon}$ is locally uniformly Lipschitz in $(0,T)$. For all $(t,z)\in\Omega$, we have
$\partial_{t}\Phi^{\varepsilon}(t,z)$ $=\int_{\mathbb{R}}\partial_{t}W^{s}(t,z)\eta_{\varepsilon}(s-1)ds$
$=\int_{\mathbb{R}}(\partial_{\tau}u(st,z)-C\lvert s-1\rvert)\eta_{\varepsilon}(s-1)ds$
$=\int_{\mathbb{R}}t^{-1}\partial_{s}u(st,z)\eta_{\varepsilon}(s-1)ds-C\int_{\mathbb{R}}\lvert s-1\rvert\eta_{\varepsilon}(s-1)ds$
$=-\int_{\mathbb{R}}t^{-1}u(st,z)\eta^{\prime}_{\varepsilon}(s-1)ds-C\int_{\mathbb{R}}\lvert s-1\rvert\eta_{\varepsilon}(s-1)ds.$
Let $J\Subset(0,S)$ be a subinterval. Note that $supp(\eta_{\varepsilon}^{\prime}(s-1))\subset(1-\varepsilon_{0},1+\varepsilon_{0})$ for some sufficiently small $0<\varepsilon_{0}<1$. It follows from the local uniform Lipschitzness of $u$ on $(0,T)$ that there exists $\kappa>0$ such that
(3.11) $\lvert u(st_{1},z)-u(st_{2},z)\rvert\leq\kappa s\lvert t_{1}-t_{2}\rvert$
for all $s\in(1-\varepsilon_{0},1+\varepsilon_{0})$ and $(t_{1},z),(t_{2},z)\in J\times\Omega$. Therefore for all $(t_{1},z)$, $(t_{2},z)\in J\times\Omega$,
$\lvert\partial_{t}\Phi^{\varepsilon}(t_{1},z)-\partial_{t}\Phi^{\varepsilon}(t_{2},z)\rvert$
$=\left|\int_{\mathbb{R}}t_{1}^{-1}u(st_{1},z)\eta_{\varepsilon}^{\prime}(s-1)ds-\int_{\mathbb{R}}t_{2}^{-1}u(st_{2},z)\eta^{\prime}_{\varepsilon}(s-1)ds\right|$
$\quad\leq\left|\int_{\mathbb{R}}(t_{1}^{-1}u(st_{1},z)-t_{1}^{-1}u(st_{2},z))\eta^{\prime}_{\varepsilon}(s-1)ds\right|$
$\quad\quad+\int_{\mathbb{R}}\lvert t_{1}^{-1}-t_{2}^{-1}\rvert\lvert u(st_{2},z)\rvert\lvert\eta^{\prime}_{\varepsilon}(s-1)\rvert ds$
$\quad\leq t_{1}^{-1}2\kappa\lvert t_{1}-t_{2}\rvert\int_{\mathbb{R}}\lvert\eta^{\prime}_{\varepsilon}(s-1)\rvert ds+\frac{\lvert t_{1}-t_{2}\rvert}{t_{1}t_{2}}\lVert u\rVert_{L^{\infty}(\Omega_{T})}\int_{\mathbb{R}}\lvert\eta^{\prime}_{\varepsilon}(s-1)\rvert ds.$
Here we used (3.11) for the last inequality. This implies that for each $\varepsilon>0$, $\partial_{t}\Phi^{\varepsilon}$ is locally uniformly Lipschitz in $(0,S)$. One can also check that $\Phi^{\varepsilon}$ is locally uniformly semi-concave in $(0,S)$.
Since $W^{s}$ satisfies (3.10), we have
$(dd^{c}W^{s}(t,\cdot))^{n}$ $\geq e^{\partial_{t}W^{s}(t,\cdot)+F(t,\cdot,W^{s})}d\mu$
$\geq e^{\partial_{t}W^{s}(t,\cdot)-\partial_{t}v(t,\cdot)+F(t,\cdot,W^{s})-F(t,\cdot,v)}(dd^{c}v(t,\cdot))^{n}$
for a.e. $t\in(0,S)$. Fix such a $t$. It follows from Lemma 3.1 that
(3.12) $(dd^{c}W^{s}(t,\cdot))\wedge(dd^{c}v(t,\cdot))^{n-1}$
$\quad\geq\exp\left(\frac{\partial_{t}W^{s}(t,\cdot)-\partial_{t}v(t,\cdot)+F(t,\cdot,W^{s})-F(t,\cdot,v)}{n}\right)(dd^{c}v(t,\cdot))^{n}.$
We show that
(3.13) $(dd^{c}\Phi^{\varepsilon}(t,\cdot))\wedge(dd^{c}v(t,\cdot))^{n-1}$
$\geq\exp\left(\frac{\partial_{t}\Phi^{\varepsilon}(t,\cdot)-\partial_{t}v(t,\cdot)+F(t,\cdot,\Phi^{\varepsilon})-F(t,\cdot,v)-O(\varepsilon)}{n}\right)(dd^{c}v(t,\cdot))^{n}.$
Indeed, let $\chi(z)$ be a positive smooth test function on $\Omega$. Then we have
$\int_{\Omega}\chi(z)(dd^{c}\Phi^{\varepsilon}(t,\cdot))\wedge(dd^{c}v(t,\cdot))^{n-1}$
$\quad=\int_{\Omega}\Phi^{\varepsilon}(t,\cdot)(dd^{c}\chi)\wedge(dd^{c}v(t,\cdot))^{n-1}$
$\quad=\int_{\Omega}\left(\int_{\mathbb{R}}W^{s}(t,\cdot)\eta_{\varepsilon}(s-1)ds\right)(dd^{c}\chi)\wedge(dd^{c}v(t,\cdot))^{n-1}.$
Let us denote
$I(s,z)=\partial_{t}W^{s}(t,z)-\partial_{t}v(t,z)+F(t,z,W^{s})-F(t,z,v).$
By the Fubini’s theorem, the last term satisfies
$\int_{\mathbb{R}}\eta_{\varepsilon}(s-1)\left(\int_{\Omega}W^{s}(t,\cdot)(dd^{c}\chi)\wedge(dd^{c}v(t,\cdot))^{n-1}\right)ds$
$\quad=\int_{\mathbb{R}}\eta_{\varepsilon}(s-1)\left(\int_{\Omega}\chi(z)(dd^{c}W^{s}(t,\cdot))\wedge(dd^{c}v(t,\cdot))^{n-1}\right)ds$
$\quad\quad\geq\int_{\mathbb{R}}\eta_{\varepsilon}(s-1)\left(\int_{\Omega}\chi(z)\exp\left(\frac{I(s,z)}{n}\right)(dd^{c}v(t,\cdot))^{n}\right)ds.$
For the inequality, we used (3.12). By the Fubini’s theorem and Jensen’s inequality, the last term satisfies
(3.14) $\int_{\Omega}\chi(z)\left(\int_{\mathbb{R}}\eta_{\varepsilon}(s-1)\exp\left(\frac{I(s,z)}{n}\right)ds\right)(dd^{c}v(t,\cdot))^{n}$
$\quad\geq\int_{\Omega}\chi(z)\exp\left(\frac{1}{n}\int_{\mathbb{R}}I(s,z)\eta_{\varepsilon}(s-1)ds\right)(dd^{c}v(t,\cdot))^{n},$
By following the last part of the proof of [GLZ21-2, Theorem 6.6] and using the assumption (3.9), one can show that
$\int_{\mathbb{R}}F(t,\cdot,W^{s})\eta_{\varepsilon}(s-1)ds\geq F(t,\cdot,\Phi^{\varepsilon})-O(\varepsilon),$
which implies that
$\int_{\mathbb{R}}I(s,z)\eta_{\varepsilon}(s-1)ds$
$\quad=\int_{\mathbb{R}}\partial_{t}(W^{s}(t,z)-v(t,z))\eta_{\varepsilon}(s-1)ds+\int_{\mathbb{R}}(F(t,z,W^{s})-F(t,z,v))\eta_{\varepsilon}(s-1)ds$
$\quad\quad\geq\partial_{t}(\Phi^{\varepsilon}(t,z)-v(t,z))+F(t,z,\Phi^{\varepsilon})-F(t,z,v)-O(\varepsilon).$
Therefore the last term of (3.14) is larger than
$\int_{\Omega}\chi(z)\exp\left(\frac{\partial_{t}(\Phi^{\varepsilon}(t,z)-v(t,z))+F(t,z,\Phi^{\varepsilon})-F(t,z,v)-O(\varepsilon)}{n}\right)(dd^{c}v(t,\cdot))^{n}.$
Since $\chi$ is chosen arbitrarily, we get (3.13).
It remains to show the assumption (b) in Lemma 3.4. To show this, we prove that
  1. (i)
    there exists a function $\eta(t)\in C([0,S))$ such that $\eta(0)=0$ and
    $v(t,z)\geq v_{0}(z)-\eta(t)$
    for $dt\wedge d\mu$-a.e. $(t,z)\in\Omega_{T}$,
  2. (ii)
    for each open $U\Subset\Omega$ and $\varepsilon>0$, there exists $A=A(\varepsilon)>0$ such that
    $\int_{U}(\Phi^{\varepsilon}(t,\cdot)-O(\varepsilon)t)d\mu\leq\int_{U}h_{1}(0,\cdot)d\mu+At$
and apply Lemma 3.6. Note that (i) follows from Lemma 3.3. To prove (ii), it suffices to show that for each open $U\Subset\Omega$, there exists $A_{0}>0$ such that
(3.15) $\int_{U}W^{s}(t,\cdot)d\mu\leq\int_{U}h_{1}(0,\cdot)d\mu+A_{0}t.$
If so, by integrating each side with respect to $\eta_{\varepsilon}(s-1)ds$, we get
$\int_{U}\Phi^{\varepsilon}(t,\cdot)d\mu\leq\int_{\mathbb{R}}\left(\int_{U}h_{1}(0,\cdot)d\mu+A_{0}t\right)\eta_{\varepsilon}(s-1)ds=\int_{U}h_{1}(0,\cdot)d\mu+A_{0}t,$
which implies (ii). The proof of (3.15) is almost same with the proof of [GLZ21-2, Lemma 3.10]. Indeed, fix an open subset $U\Subset\Omega$ and let $\chi(z)$ be a positive test function on $\Omega$. Recall that $W^{s}$ satisfies
$dt\wedge(dd^{c}W^{s})^{n}\geq e^{\partial_{t}W^{s}+F(t,z,W^{s})}dt\wedge d\mu.$
It follows from the elliptic Chern-Levine-Nirenberg inequality that there exists a constant $C>0$ satisfying
$\int_{U}\chi(z)e^{\partial_{t}W^{s}(t,\cdot)+F(t,\cdot,W^{s})}d\mu\leq\int_{U}\chi(z)(dd^{c}W^{s}(t,\cdot))^{n}\leq C$
for a.e. $t\in(0,T)$. By applying Jensen’s inequality, there exists $A_{0}>0$ satisfying
$\int_{U}\chi(z)\partial_{t}W^{s}(t,\cdot)d\mu\leq A_{0},$
which implies that
$\int_{U}W^{s}(t,\cdot)d\mu\leq\int_{U}h_{1}(0,\cdot)d\mu+A_{0}t.$
Hence it follows from Lemma 3.4 and Lemma 3.6 that $\Phi^{\varepsilon}-O(\varepsilon)t\leq v$ in $\Omega_{S}$. Since $\Phi^{\varepsilon}\rightarrow u$ pointwisely as $\varepsilon\rightarrow 0$, we get the result.
Finally, we remove our assumption (3.9). Let $\kappa_{F}>0$ be a constant satisfying
$\lvert F(t_{1},z,u)-F(t_{2},z,u)\rvert\leq\kappa_{F}\lvert t_{1}-t_{2}\rvert$
for any $t_{1},t_{2}\in[0,S]$. Fix $0<\varepsilon_{0}<S-T$ and let $U^{\varepsilon_{0}}(t,z):=u(t+\varepsilon_{0},z)-\kappa_{F}\varepsilon_{0}t$. Then $U^{\varepsilon_{0}}$ satisfies
$dt\wedge(dd^{c}U^{\varepsilon_{0}}(t,z))^{n}$ $=dt\wedge(dd^{c}u(t+\varepsilon_{0},z))^{n}$
$\quad\geq e^{\partial_{t}u(t+\varepsilon_{0},z)+F(t+\varepsilon_{0},z,u(t+\varepsilon_{0},z))}dt\wedge d\mu$
$\quad\geq e^{\partial_{t}U^{\varepsilon_{0}}+\kappa_{F}\varepsilon_{0}+F(t+\varepsilon_{0},z,U^{\varepsilon_{0}})}dt\wedge d\mu$
$\quad\geq e^{\partial_{t}U^{\varepsilon_{0}}+F(t,z,U^{\varepsilon_{0}})}dt\wedge d\mu.$
For the second inequality, we used the assumption that $F$ is increasing in the third variable. Moreover, $U^{\varepsilon_{0}}$ is locally uniformly semi-concave in $(0,S)$ with the Cauchy-Dirichlet boundary data $h_{1}^{\varepsilon_{0}}:=h_{1}-\kappa_{F}\varepsilon_{0}t$. Hence one may replace $u$ by $U^{\varepsilon_{0}}$ to conclude that $U^{\varepsilon_{0}}\leq v$ in $\Omega_{S}$. Since $U^{\varepsilon_{0}}\rightarrow u$ pointwisely as $\varepsilon_{0}\rightarrow 0$ we get the result. ∎
From this theorem, we get the following uniqueness result.
Corollary 3.9.
Let $u_{0}\in PSH(\Omega)\cap L^{\infty}(\Omega)$ and $h\in C([0,T)\times\partial\Omega)$ satisfy
  • $\lim_{z\rightarrow\partial\Omega}u_{0}(z)=h(0,\cdot)$,
  • $h$ satisfies (1.3) and (1.4).
Then there exists a unique $u\in\mathcal{P}(\Omega_{T})\cap L^{\infty}(\Omega_{T})$ which is locally uniformly semi-concave in $(0,T)$ and satisfy
$\begin{cases}&dt\wedge(dd^{c}u)^{n}=e^{\partial_{t}u+F(t,z,u)}dt\wedge d\mu\text{ in }\Omega_{T},\\ &\lim_{(t,z)\rightarrow(\tau,\zeta)}u(t,z)=h(\tau,\zeta)\text{ for all }(\tau,\zeta)\in[0,T)\times\partial\Omega,\\ &\lim_{t\rightarrow 0+}u(t,\cdot)=h(0,\cdot)\text{ in }L^{1}(d\mu).\end{cases}$
Proof.
The existence of the solution comes from [Ka25, Theorem 4.5]. By Theorem 3.8, we get the uniqueness of such a solution. ∎
Remark 3.10.
Even though [Ka25, Theorem 4.5] only considered the case when $T<+\infty$, Corollary 3.9 covers the case when $T=+\infty$. Indeed, due to the uniqueness, one can obtain the solution by gluing each solutions defined on $\Omega_{T^{\prime}}$ for each $T^{\prime}<T$.
Next theorem is another application of Theorem 3.8, which explains the long-term behavior of the solution. The motivation of the proof comes from [GLZ20, Theorem 5.3].
Theorem 3.11.
Let us denote $T=+\infty$. Let $u\in\mathcal{P}(\Omega_{T})\cap L^{\infty}(\Omega_{T})$ be the unique function which is locally uniformly semi-concave in $(0,T)$ and satisfy
$\begin{cases}&dt\wedge(dd^{c}u)^{n}=e^{\partial_{t}u+F(z,u)}dt\wedge d\mu\text{ in }\Omega_{T},\\ &\lim_{(t,z)\rightarrow(\tau,\zeta)}u(t,z)=0\text{ for all }(\tau,\zeta)\in[0,T)\times\partial\Omega,\\ &\lim_{t\rightarrow 0+}u(t,\cdot)=\psi(\cdot)\text{ in }L^{1}(d\mu),\end{cases}$
where $\psi$ is given in (1.6). If there exists a constant $L_{F}>0$ such that
(3.16) $\lvert F(z,r_{1})-F(z,r_{2})\rvert\geq L_{F}\lvert r_{1}-r_{2}\rvert$
for all $z\in\Omega$ and $r_{1},r_{2}\in J$, then $u(t,\cdot)\rightarrow\psi_{\infty}$ uniformly as $t\rightarrow+\infty$ for $\psi_{\infty}\in PSH(\Omega)\cap L^{\infty}(\Omega)$ satisfying
(3.17) $\begin{cases}&(dd^{c}\psi_{\infty})^{n}=e^{F(z,\psi_{\infty})}d\mu\text{ in }\Omega,\\ &\lim_{z\rightarrow\partial\Omega}\psi_{\infty}(z)=0.\end{cases}$
Remark 3.12.
It follows from [Ka25, Section 5] that $u$ actually satisfies $\lim_{t\rightarrow 0+}u(t,z)=\psi(z)$ for all $z\in\Omega$, since $(\mu,\psi)$ is admissible (see [Ka25, Definition 5.2]).
Proof.
It follows from [Ko00, Theorem 1.1] that there exists $\psi_{\infty}$ satisfying (3.17). It remains to show that $u(t,\cdot)\rightarrow\psi_{\infty}$ as $t\rightarrow+\infty$.
Let us define
$\Phi(t):=\frac{n}{L_{F}}(e^{L_{F}t}-1)\log(e^{L_{F}t}-1)-nte^{L_{F}t}.$
One can check that $\Phi$ satisfies
  1. (i)
    $\Phi(t)\leq 0$ for all $t\in[0,T)$,
  2. (ii)
    $\Phi^{\prime}(t)=ne^{L_{F}t}\log(1-e^{-L_{F}t})$,
  3. (iii)
    $\Phi(t)e^{-L_{F}t}$ is bounded and $\Phi(t)e^{-L_{F}t}\rightarrow 0$ as $t\rightarrow\infty$.
Next, by the local uniform Lipschitzness of $F$, there exists $\kappa_{F}>0$ such that
(3.18) $\lvert F(z,r_{1})-F(z,r_{2})\rvert\leq\kappa_{F}\lvert r_{1}-r_{2}\rvert$
for all $r_{1},r_{2}\in(-M,M)$ where
$M:=\lVert\psi\rVert_{L^{\infty}(\Omega)}+\lVert\psi_{\infty}\rVert_{L^{\infty}(\Omega)}+\sup_{t\geq 0}\lvert\Phi(t)e^{-L_{F}t}\rvert$.
Let us define
$W(t,z):=e^{-L_{F}t}\psi(z)+(1-e^{-\kappa_{F}t})\psi_{\infty}(z)+\Phi(t)e^{-L_{F}t}.$
Since $\psi,\psi_{\infty}\leq 0$ on $\Omega$ and $\Phi\leq 0$ on $[0,T)$,
$\partial_{t}W+F(z,W)$ $=-L_{F}e^{-L_{F}t}(\psi+\Phi)+\kappa_{F}e^{-\kappa_{F}t}\psi_{\infty}+\Phi^{\prime}e^{-L_{F}t}+F(z,W)$
$\quad\leq F(z,\psi_{\infty}+e^{-L_{F}t}(\psi+\Phi))-L_{F}e^{-L_{F}t}(\psi+\Phi)+\Phi^{\prime}e^{-L_{F}t}$
$\quad\leq F(z,\psi_{\infty})+n\log(1-e^{-L_{F}t}).$
Here we used (3.18) for the first inequality and (3.16) for the second inequality. Therefore $W\in\mathcal{P}(\Omega_{T})\cap L^{\infty}(\Omega_{T})$ satisfies
$dt\wedge(dd^{c}W)^{n}$ $\geq(1-e^{-\kappa_{F}t})^{n}dt\wedge(dd^{c}\psi_{\infty})^{n}$
$\geq e^{n\log(1-e^{-L_{F}t})+F(z,\psi_{\infty})}dt\wedge d\mu$
$\geq e^{\partial_{t}W+F(z,W)}dt\wedge d\mu.$
Moreover, $W$ is locally uniformly semi-concave in $(0,T)$ and satisfies
$\begin{cases}&\lim_{(t,z)\rightarrow(\tau,\zeta)}W(t,z)=\Phi(\tau)e^{-L_{F}\tau}\leq 0\text{ for all }(\tau,\zeta)\in[0,T)\times\partial\Omega,\\ &\lim_{t\rightarrow 0+}W(t,z)=\psi(z)\text{ for all }z\in\Omega.\end{cases}$
Hence it follows from Theorem 3.8 that
$W(t,z)\leq u(t,z),$
which implies that
(3.19) $\psi_{\infty}(z)-u(t,z)\leq e^{-\kappa_{F}t}\psi_{\infty}(z)-e^{-L_{F}t}(\psi(z)+\Phi(t)).$
Next, let us define
$V(t,z):=(1-e^{-ct})\psi_{\infty}+e^{-ct}$
where
$c:=\frac{n}{1+\lVert\psi_{\infty}\rVert_{L^{\infty}(\Omega)}}>0.$
Note that $V\in\mathcal{P}(\Omega_{T})\cap L^{\infty}(\Omega_{T})$ is locally uniformly semi-concave in $(0,T)$ and satisfies
$\begin{cases}&\lim_{(t,z)\rightarrow(\tau,\zeta)}V(t,z)=e^{-c\tau}>0\text{ for all }(\tau,\zeta)\in[0,T)\times\partial\Omega,\\ &\lim_{t\rightarrow 0+}V(t,z)=1>\psi(z)\text{ for all }z\in\Omega.\end{cases}$
Moreover, since $F=F(z,r)$ is increasing in $r$,
$\partial_{t}V+F(z,V)$ $\geq ce^{-ct}(\psi_{\infty}-1)+F(z,\psi_{\infty})$
$\geq-ne^{-ct}+F(z,\psi_{\infty})$
$\geq n\log(1-e^{-ct})+F(z,\psi_{\infty}).$
Here we used the inequality that $V\geq\psi_{\infty}$ for the first inequality. Therefore $V$ satisfies
$dt\wedge(dd^{c}V)^{n}$ $=(1-e^{-ct})^{n}dt\wedge(dd^{c}\psi_{\infty})^{n}$
$=e^{n\log(1-e^{-ct})+F(z,\psi_{\infty})}dt\wedge d\mu$
$\quad\leq e^{\partial_{t}V+F(z,V)}dt\wedge d\mu.$
It follows from Theorem 3.8 that
$V(t,z)\geq u(t,z),$
which implies that
(3.20) $u(t,z)-\psi_{\infty}(z)\leq e^{-ct}(1-\psi_{\infty}).$
Finally, by (3.19) and (3.20), $u(t,\cdot)\rightarrow\psi_{\infty}$ uniformly as $t\rightarrow\infty$. ∎